OpenStudy (anonymous):

Can some one please help me ASAP ! ? What is the slope of the line between (-4,4) and (-1,2) a.1 b.2 c.-2 d.-1

OpenStudy (anonymous):

So,the slope is the change in y over the change in x. Pick a point

OpenStudy (anonymous):

(2-4/-1+4) m= -2/3

OpenStudy (jenniferkeryakous):

You do: m=rise/run and plug the numbers in and solve

OpenStudy (aum):

None of the answers match the slope. Is the first point (-4,-4) instead of (-4,4)? Double check each coordinate to make sure the signs and values are correct.